OpenStudy (gtxmuqsit):

If a force of 5N is exerted on a body for 2 seconds and the body cover a distance of 1 m, then how much is the power exerted on it? a)3 watt b)7 watt 3)4 watt and please do solution for me it's simple but confusing me. @sarah786

OpenStudy (cggurumanjunath):

work=force*distance=5*1 power=work/time=(5*1)/2=2.5W==3W

OpenStudy (cggurumanjunath):

@GTXMUQSIT

OpenStudy (gtxmuqsit):

thanks @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

wc.